OpenStudy (anonymous):

Let f(x) = –4x + 7 and g(x) = 10x – 6, find f(g(x)).

OpenStudy (anonymous):

\[f(g(x))=-4(10x-6)+7\] Use FOIL \[=-40x+24+7\] \[=-40x+31\]